The period of oscillation of an object in an ideal spring and mass system is 0.50 s and the amplitude is 5 cm. Determine the velocity of the body at the equilibrium point


1
Expert's answer
2021-04-23T07:26:44-0400

Solution.

"T=0.50s;"

"x_{max}=5cm=0.05m;"

"v_{max}-?;"

"v_{max}=\omega x_{max};"

"\omega = \dfrac{2\pi}{T};"

"\omega=\dfrac{2\sdot3.14 }{0.50s}=12.56s^{-1};"

"v_{max}=12.56s^{-1}\sdot0.05m=0.628m/s;"


Answer: "v_{max}=0.628m\/s."